2: How many milliliters of an aqueous solution of 0.219 M chromium(III) sulfate is needed to obtain 3.70 grams of the salt?

____ml
Answer #1

Molarity = \frac{w\times1000}{M\times V}

w is mass of solute ( chromium sulfate) = 3.70 g

M is molar mass of chromium sulfate = 392 g/mol

V is volume of solution (mL)

Given ,

Molarity = 0.219 M

Hence,

0.219 = \frac{3.70\times1000}{392\times V}

Or, V = \frac{3.70\times1000}{392\times 0.219}

Or, V = 43.1 mL.

Hence, 43.1 mL of an aqueous solution is needed.